# Controllable Logical Hypothesis Generation for Abductive Reasoning in Knowledge Graphs

### v1

@@ 4.2 Experiment Results and Analysis
We evaluated the quality and controllability of generated hypotheses on three datasets under five conditions: pattern, relation-number, entity-number, specific-entity, and specific-relation (see Section 3.3). We also report results without any control condition. Detailed results are shown in Table 1.
Compared to the unconditional setting, the model shows notable improvements in semantic quality under conditional constraints, likely due to the additional guidance that conditions provide. Across all condition types, the model consistently achieves high condition-adherence accuracy, with most values exceeding 80%. Structural constraints generally result in better semantic similarity than semantic-focused conditions, possibly because they help eliminate more irrelevant hypotheses. Among structural conditions, using a fixed format yields the best performance in both semantic and adherence metrics, as it simultaneously constrains entity count, relation count, and logical pattern. While both specific-entity and specific-relation significantly enhance semantic similarity to a similar extent, the model shows a clear preference in condition adherence toward specific-relation.
@@ 4.3 Ablation Study
We further analyzed the influence of two proposed components of CtrlHGen, dataset augmentation based on sub-logical decomposition and the reward function.
@@
Reward Function. We investigate different reward functions on WN18RR with the "pattern" condition. The results has been shown in Table 2. Reinforcement learning notably improves generalization and reduces accuracy variance compared to supervised learning. Removing Dice and Overlap rewards weakens performance, indicating that Jaccard alone is too strict and may hinder convergence. Excluding the condition-adherence reward slightly improves semantic similarity but harms condition adherence, confirming our reward design effectively balances both objectives.
@@
Figure 5: Results of ablation studies for the sub-logical decomposition
@@ 4.5 Case study
To further illustrate controllable hypothesis generation, we present two representative cases from FB15k-237. In the first, the model generates fine-grained hypotheses that cover more relevant artists by controlling the logical structure. In the second, it focuses on specific entities, uncovering both subtle and strong associations between observation and target entities. Even for seemingly unrelated ones (e.g., BAFTA Award), the model uses logical disjunctions to preserve semantic relevance and constraint compliance. The corresponding figure and detailed analysis appear in Appendix C.
@@ Appendix B More experiment Details
In this work, we constrain both observations and hypotheses to be composed solely of entities, relations from the knowledge graph, and logical connectives. To prevent semantic leakage from pre-trained language models, we represent entities and relations using their corresponding IDs in the knowledge graph. As a result, the task is essentially a sequence generation or graph generation problem, rather than a traditional natural language processing task. Accordingly, we adopt the autoregressive GPT-2 architecture as our model, and do not employ more advanced GPT variants, even though they are permissible.
@@ Appendix C More results
C.1 Case study
To further illustrate controllable hypothesis generation, we present two representative cases from FB15k-237. In the first case shown in Fig. 7, the observation is four different types of music { Blues, Jazz, Rhythm_and_blues, Bebop}. By incorporating more complex logical patterns as conditions, the model is able to generate increasingly fine-grained answers. For example, under the simplest "1p" logical pattern, the model can identify their common parent music genre. As the complexity of the logical pattern increases, the model becomes capable of answering more detailed queries, such as identifying artists associated with those music genres. In the second case shown in Fig. 8, it focuses on specific entities. For strongly related entities such as Yahoo, the model is able to identify clear connections with the observation set. Even for entities with weaker relationships, such as two movies, the model can still capture hidden associations between them. Surpringsingly, for seemingly unrelated entities like BAFTA_Award_for_Best_Sound, the model is able to generate high-quality hypotheses by leveraging a logical structure involving the "or" operator, while still adhering to the given constraints.
@@ C.1 Case study
Observation: Avid_Technology, Intel_Corporation, IBM, Microsoft_Corporation, Seattle_Mariners, Walmart, Cisco_Systems, Inc., Golden_State_Warriors, Seattle_Supersonics, Los_Angeles_Kings, San_Jose_Sharks, Pixar, Amazon.com, eBay Condition 1:  Entity contains Smokin_Aces Hypothesis 1:  $H=V_{?}:Employ(V_{?},P_{?})\land Film\_crew\_role(Smokin\_Aces,P_{?})$ Interpretations 1:  The organization/company that employs people who have served as a crew member in the film Smokin_Aces. Conclusion 1:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0   Condition 2:  Entity contains The_Phantom
@@
Conclusion 2:  Blues, Jazz, Rhythm_and_blues, Bebop.
@@
  Condition 3:  Entity contains Yahoo Hypothesis 3:  $H=V_{?}:Employ(V_{?},P_{?})\land\neg Employed(P_{?},Yahoo)\land Employed(P_{?},Avid\_Technology)$ Interpretations 3:  The organization/company that employs people who have never been employed by Yahoo, but have been employed by Avid_Technology. Conclusion 3:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0   Condition 4:  Entity contains BAFTA_Award_for_Best_Sound Hypothesis 4:  $H=V_{?}:Employ(V_{?},P_{?})\land Nominated_{f}or(P_{?},BAFTA\_Award\_for\_Best\_Sound)\lor Employed(P_{?},Los\_Angeles\_Kings)$ Interpretations 4:  The organization/company that employs people who have been for BAFTA_Award_for_Best_Sound or have been employed by Los_Angeles_Kings. Conclusion 4:  Blues, Jazz, Rhythm_and_blues, Bebop. Jaccard Score:  1.0
@@
Appendix D Limitation
Our method requires retraining on each individual knowledge graph. As a result, a separate model must be trained for each specific domain, leading to relatively high transfer costs. Additionally, we only explore the model’s performance under single-condition control and do not conduct experiments with composite conditions. However, we believe that our framework is theoretically capable of solving such tasks.
Appendix E Broader Impact
This paper aims to promote the development of abductive reasoning in knowledge graphs. Our work has contributed to the field of deep learning and has brought about many potential social consequences. It may play an important role in fields such as clinical diagnosis or scientific discovery. However, we believe there are no negative impacts that need clarification.

### v2

@@ 4.2 Experiment Results and Analysis
We evaluated the quality and controllability of generated hypotheses on three datasets under five conditions: pattern, relation-number, entity-number, specific-entity, and specific-relation (see Section 3.3). As baselines, we use AbductiveKGR (Bai et al., 2024b) under unconditional settings (denoted as uncondition) to highlight the improvements of our approach. The results are reported in Table 1. We further compare several advanced LLMs, including GPT-4o Achiam et al. (2023), Kimi K2 (Team et al., 2025), Grok-3 (xAI, 2025), and Deepseek-V3 (Liu et al., 2024a), on FB15k237 dataset under five conditions. For these models, 2-hop subgraphs of observation entities in triple form are included as part of the prompt to compensate for their lack of KG structural knowledge. For all LLMs above, we did not use the thinking mode. And their temperatures are uniformly set to 0.0. In addition, we also added one of the most advanced reasoning models, GPT5, and adopted the thinking mode. At the same time, we constructed an attempt to combine the raw model DeepSeek-V3 with RAG. Average results across five conditions are reported in Table 2, with details provided in Appendix C.1.
Compared to AbductiveKGR (uncondition), our model shows notable improvements in semantic similarity under conditional constraints, with most condition-adherence accuracies exceeding 80%. This improvement likely stems from the additional guidance provided by the control conditions (we further provide a case in Section 4.5 whether the model can handle irrelevant control conditions). Structural conditions generally outperform semantic-focused ones in semantic similarity, with the fixed-format pattern condition achieving the best results. While both specific-entity and specific-relation conditions similarly enhance semantic similarity, the model shows a clear adherence preference for specific-relation.
On the other hand, the performance of LLMs remains very poor, even on common-sense knowledge graphs. We attribute this issue to two main factors. First, LLMs lack the ability to fully comprehend structured data, while this task requires generating correct structured query graphs rather than merely capturing semantic meaning. Moreover, when the number of observed entities is large, their two-hop subgraphs expand rapidly, producing lengthy textual representations that further challenge the model. Second, the knowledge embedded in LLMs may conflict with that of the knowledge graph. For example, given an observation set containing several singers including Justin Bieber and Kendrick Lamar, Grok-3 classified them as singers who have made hip-hop music, whereas in the knowledge graph, Justin Bieber is not a hip-hop singer. Such contradictions can significantly affect performance on certain domain-specific data. For more analysis, please refer to Appendix C.1.
@@ 4.3 Ablation Study
We studied the influence of two proposed components of CtrlHGen, dataset augmentation based on sub-logical decomposition and the reward function.
@@
Reward Function. We investigate different reward functions on WN18RR with the "pattern" condition. The results has been shown in Table 3. Reinforcement learning notably improves generalization and reduces accuracy variance compared to supervised learning. Removing Dice and Overlap rewards weakens performance, indicating that Jaccard alone is too strict and may hinder convergence. Excluding the condition-adherence reward slightly improves semantic similarity but harms condition adherence, confirming our reward design effectively balances both objectives. We further analyzed the possible reasons why semantic similarity slightly decreased when conditional adherence was introduced in Appendix C.
@@
Figure 5: Results of ablation studies for the sub-logical decomposition.
@@ 4.5 Case Study
To demonstrate our controlbility we present two representative cases from FB15k-237, with results provided in Appendix. In the first case (Fig. 8), the observation consists of four music genres: {Blues, Jazz, Rhythm_and_Blues, Bebop}. As the logical pattern conditions grow in complexity, the model produces increasingly fine-grained answers. For instance, under the basic “1p” pattern it identifies their common parent genre, while more complex patterns enable it to retrieve finer details such as artists associated with these genres. In the second case shown in Fig. 9, it focuses on specific entities. For strongly related entities such as Yahoo, the model is able to identify clear connections with the observation set. Even for entities with weaker relationships, such as two movies, the model can still capture hidden associations between them. Surpringsingly, for seemingly unrelated entities like BAFTA_Award_for_Best_Sound, the model is able to generate high-semantic-quality hypotheses by leveraging the logical "or" operator, while still ensuring adherence to the given constraints.
@@ 5 Conclusion
6 Acknowledgements
The corresponding author is Yangqiu Song. We owe sincerely thanks to all authors for their valuable efforts and contributions. The authors of this paper are supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China and the National Natural Science Foundation of China (No.62462007 and No.62302023).
@@ Appendix B More experiment Details
B.1 Smatch
Smatch (Cai and Knight, 2013) is an evaluation metric for Abstract Meaning Representation (AMR) graphs, which are directed acyclic graphs with two node types (variable and concept) and three edge types (instance, attribute, and relation). Given a predicted graph $G_{p}$ and a gold graph $G_{g}$, Smatch($G_{p},G_{g}$) is computed by finding an approximately optimal mapping between the variable nodes of the two graphs and matching their edges. Following the settings of  Bai et al. (2024b), we transform the hypothesis graph $G(H)$ into an AMR graph $GA(H)$ by adding virtual nodes and instance edges, and then calculate Smatch. In short, Smatch is used to measure the degree of similarity between the generated hypothesis and the ground truth in the test set.
@@ Appendix C More results
C.1 Detailed Results
Here, we reported our detailed results of LLMs’ performance in Table 4. We also showed our prompts in Fig 7. We found that large language models are sometimes greatly influenced by semantics, thus neglecting the role of correct structure. For example, when the observation is $\{\text{Librarian, Lawyer, Mathematician, Physicist, Scientist-GB}\}$, Grok-3 will answer whether they are working in a library or in a law-related profession. However, the correct query for reference is the occupation of Gottfrie Wilhelm von Leibniz or the occupation of those influenced by Italo Calvino. In this example, the large language model found the most relevant semantic content but ignored that they could not meet all situations. Even more strangely, large language models sometimes include the entities within observations in the hypotheses they generate. Given an observation $\{\text{Fever, Fatigue, Headache}\}$, LLMs did not find any drugs that could treat them or diseases with these three symptoms. Instead, it included these three observed entities and predicates belonging to a certain symptom in its logical assumptions. That is, these observations are symptoms of fever, Headache and Fatigue. We believe this is because the large language model has not fully understood the structural relationship, thus confusing the contents of the input edge and the output edge.
On the one hand, we found that GPT5(Thinking) has achieved a significant performance improvement Firstly, the model can follow the control conditions in most cases. Secondly, higher semantic similarity is achieved under all five conditions. In contrast, models are more likely to generate hypotheses with higher semantic similarities under the control of semantic content than under structural control. This might be because the model itself is better at capturing based on semantics compared to structured reasoning. However, they still have a considerable gap compared to CtrlHGen, indicating that abductive reasoning tasks with structured knowledge remain challenging for advanced large language models.
On the other hand, Deepseek-V3 with RAG has improved performance under the condition of semantic control, but the results remains almost unchanged under the condition of structural control. We believe this can be attributed to two primary reasons: First, RAG primarily enhances semantic retrieval, enabling the model to fetch more semantically relevant context. It offers limited benefit when precise structural constraints are imposed, as these require strict path conformance rather than mere semantic relevance. Second, the provided 2-hop subgraph already serves as a highly informative prompt. Since the depth of all 13 predefined logical patterns is 2, this 2-hop subgraph covers most of the structural information required for hypothesis generation.
The consistently poor performance under structural control instead reveals the models’ persistent weakness in complex structural reasoning over graphs. Compared to standard KGQA, which only requires interpreting and following one given logical chain, abductive reasoning is fundamentally more challenging: it demands that the model simultaneously consider all relevant logical chains surrounding a set of observed entities and abduce the single most explanatory multi-hop hypothesis. This inverse, open-ended search process imposes significantly greater demands on structural understanding and logical synthesis, an area where current LLMs still fall short.
We also compared the experimental results of GPT5 (thinking) under different temperature settings on FB15k237 dataset under the ’patttern’ condition. The results are shown in Table 5. We found that a temperature of 0.0 can ensure a balance between semantic similarity and condition adherence. Excessively high temperatures may enhance the ability to explore, thereby improving semantic similarity, but they will significantly reduce condition adherence.
Table 4: The results of controllable abductive reasoning under different conditions. (Result: average score $\pm$ standard deviation.)
Dataset Condition Semantic Similarity Condition Adherence
Jaccard Dice Overlap Accuracy Smatch
GPT-4o + 2-hop subgraph pattern 4.7±0.19 5.1±0.20 7.7±0.26 85.3±0.18 55.6±0.29
relation-number 1.9±0.08 2.8±0.11 5.4±0.19 74.4±0.49 44.1±0.26
entity-number 2.2±0.08 3.3±0.12 6.0±0.20 84.3±0.36 45.3±0.22
specific-entity 2.5±0.12 3.2±0.14 5.0±0.21 77.8±0.24 20.7±0.27
specific-relation 0.9±0.06 1.3±0.08 2.4±0.13 65.5±0.26 23.8±0.23
Kimi K2 + 2-hop subgraph pattern 3.1±0.10 4.6±0.14 7.7±0.22 82.4±0.33 50.2±0.21
relation-number 2.4±0.09 3.6±0.12 8.5±0.26 71.1±0.49 47.0±0.19
entity-number 2.2±0.09 3.2±0.12 6.0±0.20 62.3±0.41 35.8±0.19
specific-entity 4.2±0.18 5.7±0.20 10.8±0.26 69.0±0.30 38.4±0.25
specific-relation 3.6±0.11 5.2±0.15 9.5±0.26 73.4±0.19 40.5±0.24
Grok-3 + 2-hop subgraph pattern 3.8±0.11 5.7±0.15 12.0±0.28 83.0±0.37 61.2±0.24
relation-number 1.8±0.07 2.8±0.10 4.6±0.17 70.5±0.45 40.0±0.26
entity-number 1.9±0.07 2.8±0.11 4.9±0.18 70.9±0.45 42.2±0.23
specific-entity 2.7±0.12 3.7±0.14 6.0±0.22 76.3±0.31 38.8±0.27
specific-relation 2.3±0.08 3.4±0.11 7.2±0.22 77.2±0.32 35.4±0.27
Deepseek-V3 + 2-hop subgraph pattern 2.7±0.10 4.0±0.13 7.1±0.21 79.1±0.50 47.2±0.31
relation-number 0.9±0.04 1.3±0.06 5.5±0.22 70.6±0.31 39.1±0.32
entity-number 1.3±0.08 1.7±0.10 3.4±0.16 69.2±0.29 40.3±0.22
specific-entity 3.7±0.15 4.7±0.17 10.2±0.29 75.8±0.30 41.6±0.28
specific-relation 1.7±0.09 2.3±0.11 5.4±0.20 74.2±0.28 40.6±0.23
GPT5(Thinking)+2-hop subgraph pattern 14.8±0.30 17.4±0.32 30.6±0.42 83.8±0.37 71.5±0.31
relation-number 14.6±0.30 17.1±0.32 31.5±0.44 96.6±0.17 56.8±0.17
entity-number 17.8±0.30 22.0±0.33 44.9±0.46 95.3±0.21 54.2±0.19
specific-entity 24.1±0.36 27.4±0.39 40.1±0.49 94.2±0.26 31.9±0.21
specific-relation 22.1±0.34 25.8±0.37 39.6±0.46 94.5±0.22 28.1±0.20
Deepseek-V3 + RAG pattern 2.8±0.09 3.8±0.22 6.1±0.21 78.5±0.34 48.2±0.35
relation-number 1.6±0.08 2.3±0.11 3.8±0.19 69.3±0.40 34.5±0.26
entity-number 0.8±0.04 1.4±0.07 3.8±0.19 72.3±0.40 39.2±0.24
specific-entity 13.7±0.31 15.4±0.33 23.0±0.42 82.8±0.41 25.9±0.24
specific-relation 7.6±0.27 10.5±0.30 15.4±0.36 80.2±0.30 16.8±0.26
Input: Observation, Logic Patterns, 2-hop subgraph, condition [Format Specification]: ’answer: <hypothesis>’. [Instructions]: Do not add quotes, explanations, or extra text and only replace <hypothesis> with your generated content. [Task]: Now you need to do the abductive reasoning in knowledge graphs. The observation (consist of entity id and semantic) is <observation>. Your task is to generate the hypothesis whose form is first order logic in <logic patterns>, where i means intersection, u means union, n means negation. p denotes the relation and e is the entity. Here is the related 2-hop subgraph <2-hop subgraph> for you, which may help you. For each (u,v,k), u is the source node, v is the target node, k is the relation. Each form is ’id: Semantic content’. Now generate the hypothesis, with the format in the <logic patterns> . Please note that you need to make sure the hypothesis you generate satisfy the <condition> .
Figure 7: Prompt Example.
Table 5: Temperature sensitivity experiment.
Temperature Semantic Similarity Condition Adherence Average
Jaccard Dice Overlap Accuracy Smatch
t=1.0 15.8±0.31 18.2±0.33 28.5±0.41 75.3±0.43 68.4±0.18 41.2
t=0.5 13.4±0.28 15.8±0.31 28.8±0.41 79.2±0.40 69.1±0.18 41.2
t=0.0 14.8±0.30 17.4±0.32 30.6±0.42 83.8±0.37 71.5±0.31 43.6
C.2 Analysis for Condition Adherence Reward
We have further analyzed the ablation study presented in the paper, comparing the performance of each logical pattern with and without the Condition Adherence(CA) reward. The results are summarized in Table 6 and Table 7.
  • For logical patterns involving negation (such as 2in, pin, and inp), we observed that even without conditional adherence rewards, the model is often able to identify the correct logical structure on its own and achieve higher accuracy. In these cases, enforcing constraint adherence may overly limit the model’s exploratory flexibility, leading to suboptimal semantic performance.
  • In contrast, for logical patterns that involve only intersection (such as pi, ip, 2i, 3i), we found a strong correlation between improved constraint satisfaction and enhanced semantic similarity. Without reinforcement signals guiding the model to comply with constraints, it tends to generate alternative formats that deviate from the intended structure, resulting in decreased semantic quality.
  • Interestingly, for the ‘3in’ pattern, the model appears to strike a balance between intersection and negation: regardless of whether constraints are enforced, the resulting hypotheses exhibit comparable semantic similarity.
Table 6: Ablation Results of Jaccard Score
logical pattern 2in pin inp pni up 2u 3in 1p 2p pi ip 2i 3i
CtrlHGen(w/o RL) 63.6 68.1 67.6 65.2 67.0 81.9 69.2 75.4 79.3 72.6 73.8 70.1 74.8
CtrlHGen(w/o CA) 76.2 75.9 72.3 71.1 71.6 85.3 70.4 91.2 85.1 75.8 82.1 78.9 71.9
CtrlHGen 71.7 73.2 69.7 69.1 70.3 84.9 70.2 91.3 85.3 76.4 82.8 79.8 77.2
Difference -4.5 -2.7 -2.6 -2.0 -1.3 -0.4 -0.2 0.1 0.2 0.6 0.7 0.9 5.3
Table 7: Ablation Results of Condition Adherence Accuracy
logical pattern 2in pin inp pni up 2u 3in 1p 2p pi ip 2i 3i
CtrlHGen (w/o RL) 58.7 60.1 98.2 78.7 98.5 93.9 93.2 45.5 84.6 88.5 91.4 96.7 70.6
CtrlHGen (w/o CA) 82.4 84.7 78.7 79.1 91.6 97.0 34.2 65.9 74.8 57.0 58.4 76.9 16.5
CtrlHGen 84.5 98.6 84.4 85.8 98.7 96.3 95.4 89.0 96.4 98.2 98.3 98.6 90.1
C.3 More baselines
Here, we incorporated the data augmentation strategy proposed in Logic-Gen (Asai and Hajishirzi, 2020) as an additional baseline. We also compared it with our method CtrlHGen and AbductiveKGR without data augmentation but only by introducing conditional tokens. Since these two methods don’t employ reinforcement learning for conditional control, we report results after supervised training, ensuring a fair comparison. We conducted experiments on the DBpedia50 dataset and selected ‘pattern’ and ‘specific-relation’ respectively to represent structural control and semantic control. The results are reported in Table 8 and 9.
The experiments reveal that, while Logic-Gen’s data augmentation indeed improves the model’s overall grasp of logical patterns, it remains inferior to our sub-logic decomposition approach. We believe this is because the sub-logic decomposition forces the model to deeply understand and compose longer, more intricate logical chains step-by-step, leading to substantially stronger reasoning capability on complex hypotheses. It more effectively mitigates hypothesis space collapse, thereby significantly enhancing compliance when strict structural conditions are imposed.
Table 8: Results on DBpedia50 dataset under the ‘pattern’ condtion.
Model Semantic Similarity Condition Adherence
Jaccard Dice Overlap Accuracy Smatch
AbductiveKGR+condition token 68.2±0.34 73.2±0.32 80.6±0.29 66.6±0.47 77.5±0.20
Logic-Gen 69.5±0.34 73.5±0.32 79.9±0.30 65.9±0.47 77.5±0.21
CtrlHGen 70.1±0.33 74.0±0.31 80.8±0.29 73.1±0.41 80.8±0.17
Table 9: Results on the DBpedia50 dataset under the ‘specific-relation’ condition.
Model Semantic Similarity Condition Adherence
Jaccard Dice Overlap Accuracy Smatch
AbductiveKGR + condition token 69.3±0.35 73.0±0.33 86.4±0.31 78.6±0.40 58.0±0.23
Logic-Gen 70.4±0.35 73.4±0.33 88.0±0.29 75.9±0.42 54.9±0.23
CtrlHGen 72.7±0.33 77.2±0.31 90.7±0.27 80.0±0.40 51.6±0.23
C.4 Case study
In this section, we show the results of two case study in Fig 8 and Fig 9.
@@ C.4 Case study
Observation: Avid_Technology, Intel_Corporation, IBM, Microsoft_Corporation, Seattle_Mariners, Walmart, Cisco_Systems, Inc., Golden_State_Warriors, Seattle_Supersonics, Los_Angeles_Kings, San_Jose_Sharks, Pixar, Amazon.com, eBay Condition 1: Entity contains Smokin_Aces Hypothesis 1: $H=V_{?}:Employ(V_{?},P_{?})\land Film\_crew\_role(Smokin\_Aces,P_{?})$ Interpretations 1: The organization/company that employs people who have served as a crew member in the film Smokin_Aces. Conclusion 1: Avid_Technology, Intel_Corporation, IBM, Microsoft_Corporation, Seattle_Mariners, Walmart, Cisco_Systems, Inc., Golden_State_Warriors, Seattle_Supersonics, Los_Angeles_Kings, San_Jose_Sharks, Pixar, Amazon.com, eBay. Jaccard Score: 1.0   Condition 2: Entity contains The_Phantom
@@
Conclusion 2: Avid_Technology, Intel_Corporation, IBM, Microsoft_Corporation, Seattle_Mariners, Walmart, Cisco_Systems, Inc., Golden_State_Warriors, Seattle_Supersonics, Los_Angeles_Kings, San_Jose_Sharks, Pixar, Amazon.com, eBay.
@@
  Condition 3: Entity contains Yahoo Hypothesis 3: $H=V_{?}:Employ(V_{?},P_{?})\land\neg Employed(P_{?},Yahoo)\land Employed(P_{?},Avid\_Technology)$ Interpretations 3: The organization/company that employs people who have never been employed by Yahoo, but have been employed by Avid_Technology. Conclusion 3: Avid_Technology, Intel_Corporation, IBM, Microsoft_Corporation, Seattle_Mariners, Walmart, Cisco_Systems, Inc., Golden_State_Warriors, Seattle_Supersonics, Los_Angeles_Kings, San_Jose_Sharks, Pixar, Amazon.com, eBay. Jaccard Score: 1.0   Condition 4: Entity contains BAFTA_Award_for_Best_Sound Hypothesis 4: $H=V_{?}:Employ(V_{?},P_{?})\land Nominated\_for(P_{?},BAFTA\_Award\_for\_Best\_Sound)\\ \lor Employed(P_{?},Los\_Angeles\_Kings)$ Interpretations 4: The organization/company that employs people who have been nominated for BAFTA_Award_for_Best_Sound or have been employed by Los_Angeles_Kings. Conclusion 4: Avid_Technology, Intel_Corporation, IBM, Microsoft_Corporation, Seattle_Mariners, Walmart, Cisco_Systems, Inc., Golden_State_Warriors, Seattle_Supersonics, Los_Angeles_Kings, San_Jose_Sharks, Pixar, Amazon.com, eBay. Jaccard Score: 1.0
@@
C.5 Multi-Dialogue Case
In this section, we implemented a simple yet highly interactive multi-round dialogue system that automatically adjusted control conditions based on the user’s evolving intentions and the outcomes of previous rounds. We leveraged a large language model (DeepSeek-V3) to intelligently select appropriate control conditions according to the user’s expressed intent. The prompt used for this condition-selection LLM is presented in Fig. 10. At each turn, the LLM generated updated control conditions by jointly considering the hypothesis produced in the previous round, its derived conclusions, the corresponding Jaccard similarity score, and the current user input. These dynamically selected conditions were then passed back to the core hypothesis generation model. A complete interaction example is shown in Fig. 11.
In this case, the initial observation consisted of four songs. In the first round, the user expressed interest in connections related to the acoustic guitar. The system accordingly generated a relatively broad hypothesis that slightly over-covered the observed entities. In the second round, the user asked who the artist was; the LLM selected “specific-relation” as the control condition to focus the generation. Although a relevant hypothesis was produced, it remained somewhat vague. Consequently, in the third round, the user requested a simpler logical structure. The LLM responded by enforcing the simplest available logic pattern, successfully revealing that all four songs were authored by Tracy Lawrence. Finally, wishing to explore the observation more deeply, the user sought additional related information. The LLM then imposed a relation count of three as the control condition, prompting the model to generate a richer, more complex hypothesis that incorporated two different associated artists.
Through this multi-round interaction, the system seamlessly combines structural and semantic control signals, gradually improving the relevant hypotheses closely related to the user’s constantly evolving exploration goals. It demonstrates the potential of our method in real-world scenarios.
Input: Observation, Hypothesis(last round), Condition(last round),Jaccard_score(last round), Intention, Logic Patterns [Instructions]: Do not add quotes, explanations, or extra text and only replace <hypothesis> with your generated content. [Task]: Now I am doing the abductive reasoning in knowledge graphs. The observation (consist of entity id: semantic) is Observation. I have generated a hypothesis <Hypothesis> under the condition <Condition>. The conclusion of this hypothesis is <conclusion>. And the jaccard score between the conclusion and the observation is <Jaccard_score>. My intention is that <Intention>. Now your task is to adjust the condition and my model will generate a new hypothesis under the condition to make the jaccard score between the conclusion and the observation higher and conforms to my intention. [Format Specification]: The condition can be one of them ’pattern’, ’entitynumber’, ’entity’, ’relation’, ’relationnumber’. For the pattern, you can specify the corresponding first-order logical hypothesis format including <Logic Patterns>. If you choose this condition, your response must be condition: <pattern: <pattern_name>>. For the entitynumber, you can specify the number of entities in the hypothesis. If you choose this condition, your response must be condition: <entitynumber: <number>>. For the entity, you can specify the entities in the hypothesis. If you choose this condition, your response must be condition: <entity: <entity_name>>. For the relation, you can specify the relations in the hypothesis. If you choose this condition, your response must be condition: <relation: <relation_name>>. For the relationnumber, you can specify the number of relations in the hypothesis. If you choose this condition, your response must be condition: <relationnumber: <number>>.
Figure 10: Prompt for LLM.
Observation: Alibis_(song), I_See_It_Now_(song), If_the_Good_Die_Young, Somebody_Paints_the_Wall Intention 1: I want to know about the connection between these songs and the guitar. Condition 1: Entity contains Acoustic_guitar. Hypothesis 1: $H=V_{?}:WorkofArtist(V_{?},P_{?})\land Instrumentof(Acoustic\_guitar,P_{?})$ Interpretations 1: The songs that are the work of the artist whose instrument is Acoustic_guitar. Conclusion 1: Alibis_(song), I_See_It_Now_(song), If_the_Good_Die_Young, Somebody_Paints_the_Wall, If_I_Had_Any_Pride_Left_at_All. Jaccard Score: 0.66   Intention 2: I want to know which artist is the author of these songs.
Condition 2: Relation contains WorkofArtist.
Hypothesis 2: $H=V_{?}:WorkofArtist(V_{?},P_{?})\land AssociateArtist(Ken\_Mellons,P_{?})$
Interpretations 2: The songs that are the work of the artist whose associateArtist is Ken_Mellons.
Conclusion 2: Alibis_(song), I_See_It_Now_(song), If_the_Good_Die_Young, Somebody_Paints_the_Wall.
Jaccard Score: 1.0
  Intention 3: This is too complex. I want to make the logic simpler.
Condition 3: Logic Pattern 1p.
Hypothesis 3: $H=V_{?}:WorkofArtist(V_{?},Tracy\_Lawrence)$
Interpretations 3: The songs that are the work of the artist Tracy_Lawrence.
Conclusion 3: Alibis_(song), I_See_It_Now_(song), If_the_Good_Die_Young, Somebody_Paints_the_Wall.
Jaccard Score: 1.0
  Intention 4: I want to know more. Let’s explore more.
Condition 4: Logic contains 3 relations.
Hypothesis 4: $H=V_{?}:Musicband(V_{?},P_{?})\land AssociateArtist(Josh\_Logan\_(country\_singer),P_{?})\land AssociateArtist(Ken\_Mellons,P_{?})$
Interpretations 4: The songs that are played by the musical band whose associate artists are Josh_Logan_(country_singer) and Ken_Mellons.
Conclusion 4: Alibis_(song), I_See_It_Now_(song), If_the_Good_Die_Young, Somebody_Paints_the_Wall.
Jaccard Score: 1.0
Figure 11: Case Study of Multi-round Dialogue.
Appendix D The use of LLMs
In this paper, large language models (LLMs) were employed exclusively for language refinement, such as improving grammar, clarity, and readability of the manuscript. They were not utilized in any stage of the research process itself, including the formulation of ideas, experimental design, data collection, analysis, or interpretation of results.